Betty is 12, exactly one third as old as her father. in how many years will she be half as old as he is?Betty is 12, exactly one third as old as her father. in how many years will she be half as old as he is?
Answer by mananth(16946)

Betty is 12, exactly one third as old as her father. in how many years will she be half as old as he is?
Betty is 12 years
she is 1/3 fathers age
so father is 12*3 = 36 years
..
after x years they will be
betty = 12+x
Father = 36+x
..
12+x= 1/2 * (36+x)
multiply by 2
2(12+x)= 36+x
24+2x=36+x
-x
24+2x-x=36+x-x
24+x=36
-24
x+24-24=36-24
x=12
After 12 years she will be 1/2 father's age.
...
m.ananth@hotmail.ca